Solve:
-0.3(x-7)+x=0.5(3-x)
I tried to multiply all the terms by 100 and solve the equation and the answer I got was x=2.86 I rounded up

Answer by Alan3354(69443)

Solve:
-0.3(x-7)+x=0.5(3-x)
--------------
-3(x-7)+10x = 5(3-x)
-3x + 21 + 10x = 15 - 5x
7x + 21 = 15 - 5x
12x = -6
x = -1/2